OpenStudy (anonymous):

WILL AWARD MEDAL The population of a city is 97,434 and grows continuously at a rate of 5.2% each year. What is the approximate population of the city in 23 years?

OpenStudy (anonymous):

321,369 125,310 322,201 213,965

OpenStudy (ranga):

Use the same formula as the compound interest formula.

OpenStudy (ranga):

\[ \Large P(t) = P_0(1 + r)^t \]where \(\large P_0\) is initial population; \(\large r\) is the rate at which the population increases in decimal; \(\large t\) is the number of years; \(\large P(t)\) is the population after \(\large t\) years.

OpenStudy (anonymous):

so....P(23)=97,434(1+5.2)^23?

OpenStudy (ranga):

5.2 is in percentage but 'r' should be in decimal in the formula and so r = 0.052. P(23) = 97,434(1+0.052)^23 = 97,434(1.052)^23 = ?

OpenStudy (anonymous):

can you show me how to figure that out? i dont recall how to do this

OpenStudy (ranga):

Use your calculator or the online one to raise 1.052 to the exponent 23 and then multiply it by 97434. (Look for something like \(\Large x^y\) in the calculator. Input: 1.052 \(\Large x^y\) 23 = x 97434 =

OpenStudy (anonymous):

i keep getting this 312660.1615885902482616 :/

OpenStudy (ranga):

Yes. Since this is population the final answer should be an integer and so round it to: 312,660.

OpenStudy (anonymous):

but these are my answers 321,369 125,310 322,201 213,965

OpenStudy (ranga):

oh you have the answer choices. Let me see.

OpenStudy (ranga):

Can you go over the numbers in the original problem to make sure there are no mistakes?

OpenStudy (anonymous):

everything is correct :/

OpenStudy (ranga):

The problem is fairly straightforward. The answer is: \( \Large 97,434(1.052)^{23} = 312,660.\)

OpenStudy (anonymous):

what do you think your best educational guess would be then?

OpenStudy (ranga):

Can you post the screenshot of the original question and the answers?

OpenStudy (anonymous):

thoose are copied and pasted. but i have some more questions im stumped on if you care to help?

OpenStudy (anonymous):

those*

OpenStudy (ranga):

Well, posting a screenshot may help spot the mistake. I will have to log off in about 5 minutes.

OpenStudy (anonymous):

i dont know how to screenshot

OpenStudy (ranga):

If you see the question and answer choices on the screen, press "print screen" key located on Windows keyboard near the top right corner. That will save the image to memory. Then open a program such as Paint, paste the image (press control-V), edit out unwanted portions and upload it here. Or you can use cell phone camera.

OpenStudy (ranga):

I have to go now.

OpenStudy (ranga):

Since they are asking for the *approximate* population I will probably pick the choice closest to the correct answer of 312,660 which is A) 321,369.